# WaterMax: breaking the LLM watermark detectability-robustness-quality trade-off

### v1

@@ 3.2 Formalization
The main ingredient of our scheme is a watermark detection that outputs a $p$-value for any text input, whatever its length. This is the case for the main works in the literature, like (Aaronson & Kirchner, 2023; Kirchenbauer et al., 2023a; Christ et al., 2023) and to some extent (Kuditipudi et al., 2023) (see App. A). Of note, the computation of the $p$-value can be hazardous on texts with repetitions of token sub-sequences and Fernandez et al. (2023) indicate how to fix this issue. This means that the $p$-value $P$ computed from a random non-watermarked text is uniformly distributed over $[0,1]$ with p.d.f. $f(p|\mathcal{H}_{0})=\mathbbm{1}_{[0,1]}(p)$. For a given prompt, the generator creates $n$ independent texts, each with possibly a different number of tokens, computes their $p$-values $P_{j}\overset{\mathsf{iid}}{\sim}\mathcal{U}_{[0,1]}$, and outputs the text with the minimum $p$-value. A well-known results in statistics states that the minimum of $n$ independent $p$-values follow a beta distribution $B(1,n)$ whose p.d.f. is $f(p|\mathcal{H}_{1})=n(1-p)^{n-1}$. The detector receives an unknown text and decides between two hypotheses:
@@
$P_{D}\triangleq\mathbb{P}(P<P_{FA}|\mathcal{H}_{1})=1-(1-P_{FA})^{n}.$ (2)
We point the reader to two important advantages:
  • The performance of the test does not depend on the choice of the detection scheme.
  • The performance of the test does not depend on the length of the text. Consequently, even an extremely small text can be watermarked in theory.
Figure 2 plots the power of the test as a function of $n$ for various probabilities of false alarm. Whatever the choice of $P_{FA}$, one can clearly observe that a huge number of generated texts ($\gg 50$) is required to obtain a power greater than $0.5$, which is unacceptable in practice. Section 4 shows how to improve this base algorithm to reach arbitrarily high power with a smaller cost in terms of computational power.
Figure 2: Theoretical power of the uniformly most powerful test (2) as a function of the number of generated texts $n$ for different probability of false-alarm.
4 Watermarking chunks of text
This section devises ways to efficiently explore the space of possible texts to find a low $p$-value. The main idea is to split the generation into $N$ iterations, each iteration creating a chunk of the text. For each chunk, a batch of $n$ text drafts is generated independently based on the text candidates of the previous iteration. This allows us to reduce the computational burden by generating only small chunks of text at each iteration while exploring many possible texts. We choose a Viterbi-like algorithm among several exploration strategies ranging from greedy search to Monte Carlo Tree Search (Chaffin et al., 2022).
4.1 Exploring the text space
Ideally, one would generate $n$ drafts at each chunk for each previous candidate, resulting in a tree of $n^{N}$ texts from which to choose the lowest $p$-value. Obviously, this is not computationally feasible. We reduce the complexity by keeping only the $m$ best candidates at each step, akin to a beam-search or a Viterbi algorithm (Viterbi, 1967). This is suboptimal because the candidates minimizing the $p$-value at a given step may not be the best once completed at iteration $N$. The pseudo-code is summarized in Alg. 1.

### v2

@@ A.2 Variants
For a given prompt, the generator creates $n$ independent texts, each with possibly a different number of tokens, computes their $p$-values $P_{j}\overset{\mathsf{iid}}{\sim}\mathcal{U}_{[0,1]}$, and outputs the text with the minimum $p$-value. Note that in order for the $p$-value to follow a uniform distribution under $\mathcal{H}_{0}$, we assume the c.d.f of the scores to be continuous. A well-known results in statistics states that the minimum of $n$ independent $p$-values follow a beta distribution $B(1,n)$ whose p.d.f. is $f(p|\mathcal{H}_{1})=n(1-p)^{n-1}$.
The detector receives an unknown text and decides between two hypotheses:
@@
[…]
Appendix C Proof of proposition 4.1
The received text is composed of chunks whose local $p$-values are distributed as beta distributions $B(1,1)$ (i.e. $\mathcal{U}_{[0,1]}$) under $\mathcal{H}_{0}$ or $B(1,n)$ under $\mathcal{H}_{1}$ when there is no attack (see Sect. B). Knowing that if $X\sim B(1,\beta)$ then $-\log(1-X)\sim\mathcal{E}(\beta)$, $\Lambda(\mathbf{P})$ is the sum of independent exponential r.v. and thus follows a Gamma distribution under both hypothesis but with different parameters:
This leads to the characterization of the test’s performance:
$P_{FA}$ $=\gamma_{N,1}\left(\tau\right),$ (16)
where $\gamma_{x,y}$ is the lower incomplete gamma function with shape parameter $x$ and scale parameter $y$.
Appendix D Proof of proposition 5.1
Under $\mathcal{H}_{1}$, $\Lambda_{\mathsf{rob}}(\mathbf{U})$ can be written as the sum of $N$ independent chunk scores: $\Lambda_{\mathsf{rob}}(\mathbf{U})=\sum_{i=1}^{N}M_{i}^{(n)}$. These chunk scores are distributed as the maximum of $n$ Gaussian variables $\mathcal{N}(0;\ell)$, corresponding to the score of the drafts. Those maxima have the c.d.f. $F_{M^{(n)}}(x)=\Phi\left(x/\sqrt{\ell}\right)^{n}$ and one can compute numerically the expectation $\mathbb{E}(M^{(n)})=\sqrt{\ell}e(n)$ and the variance $\mathbb{V}(M^{(n)})=\ell v(n)$, see Tab. 1. As far as we know, there is no closed-form for these quantities, but classical results in concentration inequalities for the maximum show that $e(n)$ scales as $\sqrt{\log(n)}$ and $v(n)$ as $1/\log(n)$ [4].
$\mu_{1}$ $=$ $N\sqrt{\ell}e(n)=\sqrt{NL}e(n),$ (18)
$\sigma_{1}^{2}$ $=$ $N\ell v(n)=Lv(n).$ (19)
This gives the following power of the test:
$n$ 1 2 3 4 5 6 7 8 9 10
$e(n)$ 0 0.56 0.84 1.03 1.16 1.26 1.35 1.42 1.48 1.54
$v(n)$ 1 0.68 0.56 0.49 0.45 0.42 0.40 0.37 0.36 0.34
Appendix E Proof of proposition 5.2
$\left(\mathbf{1}_{\alpha L}^{\top}\mathbf{A}\mathbf{U}\mid\sum_{i=0}^{L}U_{i}=s\right)\sim\mathcal{N}\left(\alpha s;L\alpha(1-\alpha)\right).$ (22)
According to App. D, an approximation for $N$ large enough is $S\sim\mathcal{N}(\mu_{1};\sigma_{1}^{2})$. The law of total variance gives
Note that this analysis is general in the sense that it does not depend on the existence of an attack at a given $\alpha$ resulting in an acceptable quality. We believe the role of the watermark designer is to ensure the highest possible detectability whatever the strength of the attack; the design of attacks preserving text quality is the burden of the attacker. An analysis of the robustness of a watermark should thus be comprehensive, providing the performance of the watermark under any possible type of attacks, existing or not, and independent of text quality.
Appendix F Pseudo-code of the watermark embedding